OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + y = 8 x − 2y = 2 0 2 There is no x value as there is no solution to this system x can be any value as there are infinitely many solutions to this system

OpenStudy (amistre64):

since the equations arent scalars of each other, I dont see them being none and all

OpenStudy (amistre64):

if we scale the 1st one by 2 we can eliminate our y variables and sove for x

OpenStudy (amistre64):

9x = 18 ... and that is more solvable if i did it right in me head

OpenStudy (amistre64):

we could also do a cramer rule; if we need the practice :)

OpenStudy (amistre64):

x = (2(1)+8(2))/(1(1)+4(2)) = 18/9